A management consulting team needs a team of four to analyze the operations of a new client. The team should comprise an accountant, a production specialist, a finance specialist, and a management specialist. On its staff, the consulting firm has available six accountants, five production specialists, three finance specialists, and eight management specialists. How many different teams could be formed from the available individuals?
Answer #1

here we can select a accountant from 6 accountants by 6C1 ways

we can select a production specialist from 5 production specialists by 5C1 ways

we can select a finance specialist from 3 finance specialists by 3C1 ways

we can select a management specialist from 8 management specialists by 8C1 ways

so multiply all ways for final answer

6C1 * 5C1 * 3C1 * 8C1 = 6*5*3*8=720

so total 720 ways teams can be formed from available individuals.